THERE ARE TWO BODIES OF MASSES 100 KG AND 10000 KG SEPERATED BY A DISTANCE 1 METRE. AT WHAT DISTANCE FROM THE SMALLER BODY , INTENSITY OF GRAVITATIONAL FIELD WILL BE ZERO.

Grade:9

1 Answers

Arun
6 years ago
Dear student
 
I = GM/r^2

Gx100/r^2 = Gx10000/(1-r)^2

(1-r)^2 = 100r^2

1-r = 10r

Therefore r = 1/11m
 
Regards
Arun (askIITians forum expert)